# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v3

@@ 4. Example
$g_{0}(x):=\frac{x}{-x+6},\textup{ and }g_{1}(x):=\frac{3x+1}{5-x}.$
@@
${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{6},\frac{5}{6}\right)\textup{ and }{\bf G}(\textup{Fix}(H_{1}))=(0,1).$
Let $\widetilde{\mathcal{P}_{2}}:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Then,
$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0],\ {\bf q}=(q_{0},q_{1})\in\widetilde{\mathcal{P}_{2}}.$
Hence,
$s_{2}\left({\bf G}(y)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ y\in[-1,0],\ {\bf q}\in\widetilde{\mathcal{P}_{2}}.$ (4.1)
Since $s_{2}({\bf p}|{\bf q})=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
$V\left({\bf q};\frac{6}{5}\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|(q_{0},1-q_{0})\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|\left(\frac{4}{5},\frac{1}{5}\right)\right)\leq-\frac{6}{5}\log 2\leq-0.83.$
@@
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{2}}$.}$ (4.2)
Let $\epsilon\in(0,1/4)$. Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$. Then, by (4.2) and Lemma 3.5,
@@
We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$. Hence, (3.5) holds for $N=2$, $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$. By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$, then,
@@
Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$, $0.1\leq r_{1}\leq 0.3$ and $r_{0}+r_{1}\geq 1$. Then, $r_{0}+r_{1}\leq 1.2$. The maps $f_{0}$ and $f_{1}$ can be non-affine. Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$. Hence, $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$.
In this case, we see that $a_{0}<1$ and $b_{1}+c_{1}=0$. Hence, this case corresponds to Case 2.2. However, we do not need to apply Lemma 3.11 since we have the upper bound for the relative entropy in (4.1).
@@
$g_{0}(x):=\frac{x}{8-x},\ g_{1}(x):=\frac{5x+1}{7},\textup{ and }g_{2}(x):=\frac{6}{7-x}.$
@@
${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{8},\frac{245}{344},\frac{7}{43}\right),{\bf G}(\textup{Fix}(H_{1}))=\left(\frac{1}{7},\frac{5}{7},\frac{1}{7}\right),\textup{ and }{\bf G}(\textup{Fix}(H_{2}))=\left(\frac{5}{41},\frac{175}{246},\frac{1}{6}\right).$
@@
$\widetilde{\mathcal{P}_{3}}:=\left\{(p_{0},p_{1},p_{2})\in\overline{\mathcal{P}_{3}}\ \middle|\ p_{0},p_{2}\in\left[\frac{7}{16},\frac{1}{2}\right]\right\}.$
@@
$s_{3}\left({\bf G}(y)|{\bf q}\right)\leq-0.74,\ y\in[-1/6,0],\ {\bf q}\in\widetilde{\mathcal{P}_{3}}.$ (4.3)
For $q>0$, let $f_{q}(x):=x\log(x/q)$ for $x\in[0,q]$. Then, $f_{q}$ is increasing on $[0,q/e]$ and decreasing on $[q/e,1]$. Since ${\bf q}=(q_{0},q_{1},q_{2})\in\widetilde{\mathcal{P}_{3}}$, it holds that $q_{0},q_{2}\leq 1/2$ and $q_{1}\leq 1/8$. Therefore, for $y\in[-1/6,0]$,
$s_{3}\left({\bf G}(y)|{\bf q}\right)\leq f_{1/2}(G_{0}(y))+f_{1/8}(G_{1}(y))+f_{1/2}(G_{2}(y))$
$\leq 2f_{1/2}(1/6)+f_{1/8}(2/3)=\log 3-\frac{8}{3}\log 2\leq-0.74.$
@@
By Proposition 3.6, it holds that
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.74,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{3}}$.}$ (4.4)
Let $\epsilon\in(0,1/4)$. Let $\{f_{0},f_{1},f_{2}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1,2$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$. Then, by (4.4) and Lemma 3.5,
@@
We see that $\dfrac{0.74}{\log 20}>0.24$ and $\dfrac{\log 3}{\log 3-\log 1.1}<1.1$. Hence, (3.5) holds for $N=3$, $\epsilon=0.05$, $\epsilon_{2}=0.74$ and $\delta=0.1$. By Lemma 3.4, if $\{f_{0},f_{1},f_{2}\}$ is a contractive system such that $0.05\leq r_{i}\leq 0.95$, $i=0,1,2$, $r_{0}+r_{1}+r_{2}\leq 1.1$, and $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$, then,
@@
Now we can let $\{f_{0},f_{1},f_{2}\}$ be a compatible system such that $0.475\leq r_{0}\leq 0.5$, $0.05\leq r_{1}\leq 0.1$, and $0.475\leq r_{2}\leq 0.5$. Then, $1\leq r_{0}+r_{1}+r_{2}\leq 1.1$. The maps $f_{0}$, $f_{1}$ and $f_{2}$ can be non-affine. Since $s\leq\frac{\log 3}{\log 3-\log 1.1}\leq 1.1$, $\min\{r_{0}^{s},r_{2}^{s}\}\geq(0.475)^{1.1}\geq\frac{7}{16}$. Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$.
@@
$\widetilde{\mathcal{P}_{4}}:=\left\{(p_{0},p_{1},p_{2},p_{3})\in\overline{\mathcal{P}_{4}}\ \middle|\ p_{0},p_{3}\in\left[0,\frac{1}{8}\right]\right\}.$
@@
Since $\sum_{i=0}^{3}q_{i}=1$, $q_{1}+q_{2}=1-q_{0}-q_{3}$, and hence, $\log q_{1}+\log q_{2}\leq 2\log((1-q_{0}-q_{3})/2)$. Therefore,
@@
By Proposition 3.6, it holds that
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.38,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{4}}$.}$ (4.5)
Let $\epsilon\in(0,1/4)$. Let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1,2,3$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}+r_{3}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$. Then, by (4.5) and Lemma 3.5,
@@
We see that $\dfrac{0.38}{\log 10}>0.165$ and $1.1521.15<\dfrac{\log 4}{\log 4-\log 1.2}<$ 1.152. Hence, (3.5) holds for $N=4$, $\epsilon=0.1$, $\epsilon_{2}=0.38$ and $\delta=0.2$. By Lemma 3.4, if $\{f_{0},f_{1},f_{2},f_{3}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1,2,3$, $r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, then,
@@
Now we can let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a compatible system such that $0.1\leq r_{0}\leq 0.125$, $0.4\leq r_{1}\leq 0.475$, $0.4\leq r_{2}\leq 0.475$, $0.1\leq r_{3}\leq 0.125$. Then, $1\leq r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$. The maps $f_{0}$, $f_{1}$, $f_{2}$ and $f_{3}$ can be non-affine. Since $s\geq 1$, $\max\{r_{0}^{s},r_{3}^{s}\}\leq 1/8$ and hence $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$.

### v4

@@ 4. Examples
$g_{0}(x)\coloneqq\frac{x}{-x+6}=\frac{\frac{x}{6}}{-\frac{x}{6}+1},\textup{ and }g_{1}(x)\coloneqq\frac{3x+1}{5-x}=\frac{\frac{3}{5}x+\frac{1}{5}}{-\frac{x}{5}+1}.$
@@
${\bf G}\left(\textup{Fix}(H_{0})\right)=\left(\frac{1}{6},\frac{5}{6}\right)\textup{ and }{\bf G}(\textup{Fix}(H_{1}))=(0,1).$
Let $\widetilde{\mathcal{P}_{2}}\coloneqq\left\{(p_{0},p_{1})\in\mathcal{P}_{2}\colon p_{0}\geq 4/5\right\}$. Then,
$\left|G_{0}(y)-q_{0}\right|=\left|G_{1}(y)-q_{1}\right|\geq\frac{3}{5},\ y\in[-1,0],\ {\bf q}=(q_{0},q_{1})\in\widetilde{\mathcal{P}_{2}}$.
Since $s_{2}\left({\bf p}|{\bf q}\right)=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
$\displaystyle s_{2}\left({\bf G}(y)|{\bf q}\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|(q_{0},1-q_{0})\right)$ $\displaystyle\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|\left(\frac{4}{5},\frac{1}{5}\right)\right)$
(4.1) $\displaystyle\leq-\frac{6}{5}\log 2\leq-0.83.$
@@
(4.2) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-0.83,\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z$, \ $\displaystyle{\bf q}\in\widetilde{\mathcal{P}_{2}}$.}$
Let $\epsilon\in(0,1/4)$ and let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$, $i=0,1$. Let $s\geq 1$ be the constant such that $r_{0}^{s}+r_{1}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$. Then, by (4.2) and Lemma 3.5,
@@
By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, and $r_{0}+r_{1}\leq 1.2$, then $s\leq\dfrac{\log 2}{\log 2-\log 1.2}$.
Since we see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$, equation (3.6) holds for $N=2$, $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$.
Therefore, if $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$, then
@@
Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$, $0.1\leq r_{1}\leq 0.3$ and $r_{0}+r_{1}\geq 1$. Then, $r_{0}+r_{1}\leq 1.2$. The maps $f_{0}$ and $f_{1}$ can be non-affine maps. Since $s\leq\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$, we have $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$. Hence, $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$ and we see that $\dim_{H}\mu_{\varphi}<1$.
In this case, we see that $a_{0}<1$ and $b_{1}+c_{1}=0$. Hence, this case corresponds to Case 2.2. However, we do not need to apply Lemma 3.11 since we have the direct upper bound for the relative entropy in (4.1).
@@
$g_{0}(x)\coloneqq\frac{x}{8-x},\ g_{1}(x)\coloneqq\frac{5x+1}{7},\textup{ and }g_{2}(x)\coloneqq\frac{6}{7-x}.$
@@
${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{8},\frac{245}{344},\frac{7}{43}\right),{\bf G}(\textup{Fix}(H_{1}))=\left(\frac{1}{7},\frac{5}{7},\frac{1}{7}\right),\textup{ and }$
${\bf G}(\textup{Fix}(H_{2}))=\left(\frac{5}{41},\frac{175}{246},\frac{1}{6}\right).$
@@
$\widetilde{\mathcal{P}_{3}}\coloneqq\left\{(p_{0},p_{1},p_{2})\in\overline{\mathcal{P}_{3}}\ \middle|\ p_{0},p_{2}\in\left[\frac{7}{16},\frac{1}{2}\right]\right\}.$
@@
(4.3) $s_{3}\left({\bf G}(y)|{\bf q}\right)\leq-0.74,\ y\in[-1/6,0],\ {\bf q}\in\widetilde{\mathcal{P}_{3}}$.
For $q>0$, let $f_{q}(x)\coloneqq x\log(q/x)$ for $x\in[0,1]$. Then, $f_{q}$ is increasing on $[0,q/e]$ and decreasing on $[q/e,1]$. Furthermore, for each $x\geq 0$, $f_{q}(x)$ is increasing with respect to $q$. Since ${\bf q}=(q_{0},q_{1},q_{2})\in\widetilde{\mathcal{P}_{3}}$, it holds that $q_{0},q_{2}\leq 1/2$ and $q_{1}\leq 1/8$. We also see that $\max\{G_{0}(y),G_{2}(y)\}\leq\dfrac{1}{6}<\dfrac{1}{2e}$ and $\dfrac{1}{8e}<\dfrac{2}{3}\leq G_{1}(y)$.
Therefore, for $y\in[-1/6,0]$,
$\displaystyle s_{3}\left({\bf G}(y)|{\bf q}\right)$ $\displaystyle=f_{q_{0}}(G_{0}(y))+f_{q_{1}}(G_{1}(y))+f_{q_{2}}(G_{2}(y))$
$\displaystyle\leq f_{1/2}(G_{0}(y))+f_{1/8}(G_{1}(y))+f_{1/2}(G_{2}(y))$
$\displaystyle\leq 2f_{1/2}(1/6)+f_{1/8}(2/3)=\log 3-\frac{8}{3}\log 2\leq-0.74.$
@@
By Proposition 3.7,
(4.4) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-0.74,\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z$, \ $\displaystyle{\bf q}\in\widetilde{\mathcal{P}_{3}}$.}$
[…]
@@
By Lemma 3.4, if $\{f_{0},f_{1},f_{2}\}$ is a contractive system such that $0.05\leq r_{i}\leq 0.95$, $i=0,1,2$, and $r_{0}+r_{1}+r_{2}\leq 1.1$, then $s\leq\dfrac{\log 3}{\log 3-\log 1.1}$.
Since we see that $\dfrac{0.74}{\log 20}>0.24$ and $\dfrac{\log 3}{\log 3-\log 1.1}<1.1$, equation (3.6) holds for $N=3$, $\epsilon=0.05$, $\epsilon_{2}=0.74$ and $\delta=0.1$.
Therefore, if $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$, then
@@
Now we can let $\{f_{0},f_{1},f_{2}\}$ be a compatible system such that $0.475\leq r_{0}\leq 0.5$, $0.05\leq r_{1}\leq 0.1$, and $0.475\leq r_{2}\leq 0.5$. Then, $1\leq r_{0}+r_{1}+r_{2}\leq 1.1$. The maps $f_{0}$, $f_{1}$ and $f_{2}$ can be non-affine maps. Since $\displaystyle s\leq\frac{\log 3}{\log 3-\log 1.1}\leq 1.1$, we have $\min\{r_{0}^{s},r_{2}^{s}\}\geq(0.475)^{1.1}\geq\frac{7}{16}$. Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$ and $\dim_{H}\mu_{\varphi}<1$.
@@
$\widetilde{\mathcal{P}_{4}}\coloneqq\left\{(p_{0},p_{1},p_{2},p_{3})\in\overline{\mathcal{P}_{4}}\ \middle|\ p_{0},p_{3}\in\left[0,\frac{1}{8}\right]\right\}.$
@@
Since $\displaystyle\sum_{i=0}^{3}q_{i}=1$, $q_{1}+q_{2}=1-q_{0}-q_{3}$, and hence, $\log q_{1}+\log q_{2}\leq 2\log((1-q_{0}-q_{3})/2)$. Therefore,
@@
By Proposition 3.7,
(4.5) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-0.38,\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z$, \ $\displaystyle{\bf q}\in\widetilde{\mathcal{P}_{4}}$.}$
Let $\epsilon\in(0,1/4)$ and let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$, $i=0,1,2,3$. Let $s\geq 1$ be the constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}+r_{3}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$. Then, by (4.5) and Lemma 3.5,
@@
By Lemma 3.4, if $\{f_{0},f_{1},f_{2},f_{3}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1,2,3$, and $r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$, then $s\leq\dfrac{\log 4}{\log 4-\log 1.2}$.
Since we see that $\dfrac{0.38}{\log 10}>0.165$ and $1.15<\dfrac{\log 4}{\log 4-\log 1.2}<1.152$, equation (3.6) holds for $N=4$, $\epsilon=0.1$, $\epsilon_{2}=0.38$ and $\delta=0.2$.
Therefore, if $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, then
@@
Now we can let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a compatible system such that $0.1\leq r_{0}\leq 0.125$, $0.4\leq r_{1}\leq 0.475$, $0.4\leq r_{2}\leq 0.475$, $0.1\leq r_{3}\leq 0.125$. Then, $1\leq r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$. The maps $f_{0}$, $f_{1}$, $f_{2}$ and $f_{3}$ can be non-affine maps. Since $s\geq 1$, we have $\max\{r_{0}^{s},r_{3}^{s}\}\leq 1/8$. Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, and $\dim_{H}\mu_{\varphi}<1$.